b1az3:

Mr. Howard is buying a car for $13,786. The licensing fee is $436.29, and there is a 5% sales tax on the price of the car. Estimate the combined total of the amount he is paying for the car. $15,100 $14,800 $14,200 $15,440

jhonyy9:

,, there is a 5% sales tax on the price of the car " - how you think what mean this ?

MagnusMentula:

Combine original price with the licensing fee, get 14.222.29. Multiply by 0.05 to get 5%. Add that number to 14,222.29 to get the final answer: 14,933.4045. Round as needed.

MagnusMentula:

Now that there are answer choices, I'm going to do it again, but wait to add the licensing fee. 13786 * 0.05 = 689.3 13786 + 689.3 = 14475.3 14475.3 + 436.29 = 14911.59 ... still not close to any of the given answers.